Assessing Financial Statement Effects of Transactions and Adjustments
Selected accounts of Portage Properties, a real estate management firm, are shown below as of January 31, before any accounts have been adjusted.

Debit Credit

Prepaid Insurance
$3,240
Supplies
1,540
Office Equipment
6,240
Unearned Rent Revenue
$5,550
Salaries Expense
2,325
Rent Revenue
13,250

Portage Properties prepares monthly financial statements. Using the following information, adjust the accounts as necessary on January 31 using the financial statements effect template.
(a) Prepaid insurance represents a two-year premium paid on January 1.
(b) Supplies of $710 were still available on January 31.
(c) Office equipment is expected to last eight years (or 96 months).
(d) Earlier this month, on January 1, Portage collected $5,550 for six months' rent in advance from a tenant renting space for $925 per month.
(e) Salaries of $490 have been earned by employees but yet not recorded as of January 31.

Balance Sheet

Transaction+ Cash Asset +Non Cash Assets= Liabilities+Contributed Capital + Earned capital

  Income Statement

Revenue (minus) Exspenses = Net income
Answer #1
Balance sheet Income statement
Transaction Cash
asset
+ Non-cash
assets
= Liabilities + Contributed
capital
+ Earned
capital
Revenues - Expenses = Net
income
a. + -135 = + + -135 0 - 135 = -135
b. + -830 = + + -830 0 - 830 = -830
c. + -65 = + + -65 0 - 65 = -65
d. + = -925 + + 925 925 - 0 = 925
e. + = 490 + + -490 0 - 490 = -490
Adjusting entries:
Ref. Account titles and explanation Debit Credit
a. Insurance expense (Note:1) 135
Prepaid insurance 135
(Insurance expired )
b. Supplies expense (1540-710) 830
Supplies 830
(Supplies used)
c. Depreciation expense (6240/96) 65
Accumulated depreciation 65
(Depreciation recorded)
d. Unearned rent revenue 925
Rent revenue 925
(Rent for January recorded)
e. Salaries expense 490
Salaries payable 490
(Salaries earned but not paid)
Note:1
Prepaid insurance paid for 2 years or 24 months
Insurance expense for January=(3240/24)*1=$ 135